OpenStudy (anonymous):

Travel: It took a motorboat 1 hour longer to travel 36 miles against he current than to go 36 miles with the current. The rate of the current was 3 mph. Find the rate of the boat in calm water.

OpenStudy (anonymous):

answer is 15 and the check is easy

OpenStudy (anonymous):

the work is less so, but i can write it for you if you like

OpenStudy (anonymous):

pls? I had y+3*1 and y-3*1 but I somehow think that is wrong :)

OpenStudy (anonymous):

ok we can work it out like this

OpenStudy (anonymous):

distance is 36 rate with current is x + 3 rate against current is x - 3 and using \[T=\frac{D}{R}\] we know the time with current is \[\frac{36}{x+3}\] and the time against the current is \[\frac{36}{x-3}\]

OpenStudy (anonymous):

and we know that the time against the current is one hour more than the time with the current, so we set \[\frac{36}{x+3}+1=\frac{36}{x-3}\] and solve for x

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

ok so i was almost there. I'm following your logic

OpenStudy (anonymous):

ok good. again the time against the current is one more than the time with, so if we add 1 to the time with the current we get the time against the current. that is what the equation says. now we have to solve for x which takes a bit of algebra. i will write it

OpenStudy (anonymous):

\[\frac{36}{x+3}+1=\frac{36+x+3}{x+3}=\frac{39+x}{x+3}\] so we get \[\frac{39+x}{x+3}=\frac{36}{x-3}\]

OpenStudy (anonymous):

cross multiply to get \[(x-3)(x+39)=36(x+3)\]

OpenStudy (anonymous):

what happened to the +1? Does get ignored? I understand the 36+3 and the 39-3.

OpenStudy (anonymous):

multiply out to get \[x^2-36x-117=36x+108\] giving \[x^2-225=0\] or \[x = \sqrt{225}=15\]

OpenStudy (anonymous):

we added the one to \[\frac{36}{x+3}\][

OpenStudy (anonymous):

again \[\frac{36}{x+3}+1=\frac{36}{x+3}+\frac{x+3}{x+3}=\frac{36+x+3}{x+3}= \frac{x+39}{x+3}\]

OpenStudy (anonymous):

ok...it just clicked. :)

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

got it! TY!!!!!!

OpenStudy (anonymous):

yw. btw answer is easy to check 15+3=18 and 36/18=2 15-3=12 and 36/12=3, one hour longer

OpenStudy (anonymous):

TY! You did a great job of explaining this to me!